OpenStudy (anonymous):

How many liters of water vapor can be produced if 8.9 liters of methane gas (CH4) are combusted, if all measurements are taken at the same temperature and pressure? Show all of the work used to solve this problem. CH4 (g) + 2 O2 (g) ---> CO2 (g) + 2 H2O (g)

OpenStudy (joannablackwelder):

So, this one is a little different but it is still similar.

OpenStudy (anonymous):

Still the same formula right?

OpenStudy (joannablackwelder):

Yup, but this time you have to use it twice :)

OpenStudy (anonymous):

Can you help me get started?

OpenStudy (joannablackwelder):

Sure. First we need to use the 8.9 L of methane in the ideal gas formula to find moles of methane.

OpenStudy (joannablackwelder):

Since it says same temp and pressure we can use any temp and pressure as long as we are consistent.

OpenStudy (anonymous):

Ok so P=1 V=8.9 T=273

OpenStudy (joannablackwelder):

Right. And R?

OpenStudy (anonymous):

R=0.0821

OpenStudy (joannablackwelder):

Lookin good :)

OpenStudy (anonymous):

1*8.9=n*.0821*273

OpenStudy (anonymous):

N=2.52

OpenStudy (joannablackwelder):

That's not what I get...

OpenStudy (joannablackwelder):

8.9/(.082*273)=n

OpenStudy (anonymous):

N=.397

OpenStudy (joannablackwelder):

That's what I get :)

OpenStudy (anonymous):

so we do it once more?

OpenStudy (joannablackwelder):

Not yet. What is the unit of .397?

OpenStudy (anonymous):

CH4?

OpenStudy (joannablackwelder):

Right. 0.397 mol CH4

OpenStudy (anonymous):

ok

OpenStudy (joannablackwelder):

Can you use that to get the moles of water?

OpenStudy (anonymous):

Im lost here

OpenStudy (joannablackwelder):

You can use the molar ratio of the reaction to convert from one type of mole to another,

OpenStudy (joannablackwelder):

0.397 mol CH4 (2 mol H2O/1 mol CH4) = ?

OpenStudy (anonymous):

.1985

OpenStudy (joannablackwelder):

How did you get that?

OpenStudy (anonymous):

divided .397 by two

OpenStudy (joannablackwelder):

But the 2 is in the numerator.

OpenStudy (joannablackwelder):

When the 2 is in the numerator you multiply.

OpenStudy (joannablackwelder):

When the 2 is in the denominator, you divide.

OpenStudy (anonymous):

ohh ok so if it would be 1/2 you would divide ok i got it sorry

OpenStudy (joannablackwelder):

You got it! :)

OpenStudy (joannablackwelder):

No worries

OpenStudy (anonymous):

.794

OpenStudy (joannablackwelder):

Sweet

OpenStudy (anonymous):

so this would be our new n

OpenStudy (joannablackwelder):

Right!

OpenStudy (anonymous):

P=1 n=794 R=0.0821 T=273 and we have to find V

OpenStudy (joannablackwelder):

Yup, but don't forget the decimal again :)

OpenStudy (anonymous):

.794

OpenStudy (joannablackwelder):

there you go :)

OpenStudy (anonymous):

17.796L of H20

OpenStudy (joannablackwelder):

Good job!

OpenStudy (anonymous):

Ok i just need one more

OpenStudy (joannablackwelder):

I can try :)